# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v1

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Conditioned on the event (21), and given $\tilde{S}$ specified in that event, the SUBSET algorithm succeeds provided that, for all $\ell=1,2,\dotsc,\frac{3}{2}\eta^{-}k$, and every possible $\tilde{S}^{\prime}$ formed by removing $\ell$ items from $\tilde{S}$ and replacing them with $\ell$ non-defective items, it holds that $\tilde{S}^{\prime}$ explains fewer tests than $\tilde{S}$.
@@
In view of the decoding rule in Algorithm 1, and the fact that $\tilde{S}\in\mathcal{S}_{\rm alg}$, in order to prove the first statement, it suffices to show that the sets $\tilde{S}^{\prime}$ formed in this manner (along with $\tilde{S}$ itself) collectively cover all of $\mathcal{S}_{\rm an}$ (and hence all of $\mathcal{S}_{\rm alg}$, since $\mathcal{S}_{\rm alg}\subseteq\mathcal{S}_{\rm an}$). The range $\ell=1,2,\dotsc,\frac{3}{2}\eta^{-}k$ suffices because the Hamming distance upon performing $\ell$ swaps is $2\ell$, and $\mathcal{S}_{\rm an}$ constrains the Hamming distance to be at most $3\eta^{-}k$ (see (12)).
@@
    With probability $1-o(1)$, for any $\delta_{2}=\omega(\frac{1}{\sqrt{T}})$, the number of positive tests is at most $\frac{T}{2}(1+\delta_{2})$.
@@
    With probability $1-o(1)$, under the choice $\delta_{2}=\sqrt{\frac{16\eta^{-}k\log\frac{1}{\eta^{-}}}{T}}$, it holds for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ that the number of tests in which at least one item from $\bar{s}$ is present is at least $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$.
@@
We begin with claim (i). Since each test is positive with probability $\frac{1}{2}$ and all tests are mutually independent, the number of positive tests is distributed as $\mathrm{Bin}(T,\frac{1}{2})$. Thus, due to binomial concentration (Appendix -A), it holds that there are at most $\frac{T}{2}(1+\delta_{2})$ positive tests with probability at least $1-\exp(-\frac{1}{3}\delta_{2}^{2}T)$, which is $1-o(1)$ when $\delta_{2}=\omega(\frac{1}{\sqrt{T}})$.
@@
We now use (23) along with a union bound to obtain a counterpart that holds simultaneously for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$. The distribution on the number of tests in which $\bar{s}$ is present is the same for any $\bar{s}$, and there are at most (${k\choose(1-\eta^{-})k}={k\choose\eta^{-}k}$ sets in ${\mathcal{S}}_{\mathrm{an}}$ (recall that we implicitly condition on fixed $S=s$), so the probability of any such set being present in less than $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$ tests is upper bounded by
(${k\choose\eta^{-}k}\exp\left(-2\eta^{-}k\log\frac{1}{\eta^{-}}\right)$ (24)
@@
Lemma 6 hinges on the two specified assumptions on $\eta^{-}$ and $\theta$; we now proceed to argue that it is safe to assume both of these for the purpose of proving Theorem 2. For $\eta^{-}$, this is because we only replaced the assumed scaling from $\Omega(\cdot)$ to $\Theta(\cdot)$; however, solving SUBSET with a smaller $\eta^{-}$ immediately implies the same for larger $\eta^{-}$, since one can simply remove an arbitrary subset of items from the final estimate. For $\theta$, the restriction $\theta>\frac{1}{3}$ is valid because when $\theta\leq\frac{1}{3}$, a rate of $1$ is achievable even for exact recovery [15, 17], which implies solving SUBSET for arbitrary $\eta^{-}$. Hence, for the rest of the proof, we adopt the stricter assumptions from Lemma 6, in particular noting that $\eta^{-}=o(1)$.
Using part (ii) of Lemma 6 and the fact that $\tilde{S}\in{\mathcal{S}}_{\mathrm{an}}$, we have that $\tilde{S}$ explains at least $T(1-2^{-(1-\eta^{-})(1+o(1))})$ tests with probability $1-o(1)$ (since $\tilde{S}$ explains every test it is present in due to $\tilde{S}\subseteq S$). Hence, and using part (i) of Lemma 6, the number of positive tests not explained by $\tilde{S}$ is at most the following, with probability $1-o(1)$:
@@
Suppose that the placements of defective items are such that the high-probability events (21) and (28) hold, and consider $\tilde{S}$ from (21). Fix some $\ell\in\{1,2,\dots,\frac{3}{2}\eta^{-}k\}$ and sets ${\mathcal{P}}_{\ell}\subset\tilde{S}$ and ${\mathcal{N}}_{\ell}\subset S^{\mathsf{c}}$ with $\lvert{\mathcal{P}}_{\ell}\rvert=\lvert{\mathcal{N}}_{\ell}\rvert=\ell$. Let $\tilde{S}^{\prime}=(\tilde{S}\setminus{\mathcal{P}}_{\ell})\cup{\mathcal{N}}_{\ell}$ be the set formed by swapping out the $\ell$ defective items from ${\mathcal{P}}_{\ell}$ with the $\ell$ non-defective items from ${\mathcal{N}}_{\ell}$. Then, denoting the number of tests explained by $\tilde{S}$ (resp. $\tilde{S}^{\prime}$) as $T_{\mathrm{exp}}(\tilde{S})$ (resp. $T_{\mathrm{exp}}(\tilde{S}^{\prime})$), there exists a choice of $\delta_{1}>0$ given in Lemma 4 such that
$\mathbb{P}(T_{\mathrm{exp}}(\tilde{S})$ $\leq T_{\exp}(\tilde{S}^{\prime}))\leq\exp\left(-\Omega\left(\ell\sqrt{\log n}\cdot\log n\right)\right),$ (30)
@@
Our proof strategy consists of finding a “worst case distribution” on the number of positive tests the non-defective items in ${\mathcal{N}}_{\ell}$ get placed in, and showing that the probability of $\tilde{S}^{\prime}$ explaining more tests than $\tilde{S}$ under this distribution is small. The probability of at least one of the $\ell$ non-defective items in ${\mathcal{N}}_{\ell}$ being included in any given test is $1-\big{(}1-\frac{\log 2(1+o(1))}{k}\big{)}^{\ell}=\frac{\ell\log 2}{k}\big{(}1+o(1)\big{)}$, since $\ell=o(k)$. Thus, recalling the high probability upper bound on the number of tests left unexplained by $\tilde{S}$ given in (28), the number of newly explained tests is stochastically upper bounded by the following:333 Our analysis focuses on the positive tests that these $\ell$ non-defectives get placed in. Any negative tests they get placed in will only decrease the number of explained tests further, since by Definition 4 any non-defective in a negative test does not explain any tests. However, for our purposes it suffices to naively upper bound this amount of decrease by zero.
@@
We also note the following asymptotic simplification for $12^{\eta^{-}(1+o(1))}-$ 1 when $\eta^{-}=o(1)$:
@@
With this result established, we are now in a position to show that a union bound over all possible $({\mathcal{P}}_{\ell},{\mathcal{N}}_{\ell})$, followed by a second union bound over all the values of $\ell$, still leads to the probability of $\tilde{S}$ not explaining the most tests being small. This is formalized in the following lemma.
@@
Under the setup and notation of Lemma 7, [^4] In particular, we are again conditioning on (21) and (28), and the probability is over the randomness in the placements of non-defective items. for each $\ell\in\{1,2,\dots,\frac{3\eta^{-}k}{2}\}$, define
${\mathcal{E}}_{\ell}\coloneqq\bigcup_{{\mathcal{P}}_{\ell}\subset\tilde{S}\,:\,\lvert{\mathcal{P}}_{\ell}\rvert=\ell}\bigcup_{{\mathcal{N}}_{\ell}\subset S^{\mathsf{c}}\,:\,\lvert{\mathcal{N}}_{\ell}\rvert=\ell}\left\{T_{\exp}(\tilde{S})\leq T_{\exp}((\tilde{S}\setminus{\mathcal{P}}_{\ell})\cup{\mathcal{N}}_{\ell})\right\}$ (42)
@@
$\mathbb{P}\left(\bigcup_{\ell=1}^{\frac{3}{2}\eta^{-}k}{\mathcal{E}}_{\ell}\right)\to 0.$ (43)
@@
We begin by deriving a bound on $\mathbb{P}({\mathcal{E}}_{\ell})$ for fixed $\ell$. For the event corresponding to any single choice of $({\mathcal{P}}_{\ell},{\mathcal{N}}_{\ell})$ in (42), the associated probability is upper bounded via Lemma 7. A simple counting argument reveals that there are (${(1-\eta^{-})k\choose\ell}{n-k\choose\ell}$ such pairs, which can be further upper bounded by (${k\choose\ell}{n\choose\ell}$. Hence, the union bound gives
$\mathbb{P}({\mathcal{E}}_{\ell})$ ($\leq{k\choose\ell}{n\choose\ell}\mathbb{P}\left(E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}\right)$ (44)
($\leq{k\choose\ell}{n\choose\ell}\exp\left(-\Omega\big{(}\ell\sqrt{\log n}\cdot\log n\big{)}\right)$ (45)
@@
where (46) follows since (${n\choose k}\leq n^{k}$, and (47) follows since $\ell\log k\leq\ell\log n=o(\ell\sqrt{\log n}\cdot\log n)$.
@@
$\mathbb{P}\left(\bigcup_{\ell=1}^{\frac{3}{2}\eta^{-}k}{\mathcal{E}}_{\ell}\right)$ $\leq\sum_{\ell=1}^{\frac{3}{2}\eta^{-}k}\mathbb{P}({\mathcal{E}}_{\ell})$ (48)
$\leq\sum_{\ell=1}^{\frac{3}{2}\eta^{-}k}\exp\left(-\Omega\big{(}\ell\sqrt{\log n}\cdot\log n\big{)}\right)$ (49)
@@ III-B Proof of Corollary 1
[12, Theorem 1] Fix $\alpha\in(0,1)$ and $\varepsilon>0$, and suppose that $T\geq(1-\alpha)\big{(}k\log_{2}\frac{n}{k}\big{)}(1+\varepsilon)$. Then there exists a group testing algorithm returning an estimate $\hat{S}^{\prime}$ such that $\max\{\lvert S\setminus\hat{S}^{\prime}\rvert,\lvert\hat{S}^{\prime}\setminus S\rvert\leq\alpha k$ with probability $1-o(1)$.
@@
We can use the same idea for the SUBSET problem; since the reasoning is all analogous to that of [12], we omit some details and focus mainly on the differences. We delete a fraction of items arbitrarily close to $\alpha$ and mark them as non-defective. Among those remaining, we apply the approximate recovery algorithm from [15] with $\beta=\frac{\eta^{-}}{2}$ for some $\eta^{-}=o(1)$ to obtain $\hat{S}^{\prime}$. As noted in [12], a result in [28, Appendix B] shows that this algorithm still succeeds with high probability even when the number of defectives is only known within a factor of $1\pm o(1)$. We then apply the SUBSET algorithm (Algorithm 1) using the reduced ground set and the two-sided estimate $\hat{S}^{\prime}$. While we are unable to directly apply Algorithm 1 and Theorem 2, since these assume that the number of defectives is known, we can easily modify the algorithm and proof to show the result still holds when it is only known to within a factor of $1\pm o(1)$. Specifically, letting $K^{\prime}$ be the (random) number of non-deleted defectives and defining $k^{\prime}\coloneqq\mathbb{E}K^{\prime}=(1-\alpha+\xi)k$, we observe the following:
  • A simple concentration argument reveals that $\underline{k}^{\prime}\leq K^{\prime}\leq\overline{k}^{\prime}$ with probability $1-o(1)$, where $\underline{k}^{\prime}=k^{\prime}(1-o(1))$ and $\overline{k}^{\prime}=k^{\prime}(1+o(1))$.
@@
  • Due to the above modifications, in the definition of ${\mathcal{S}}_{\rm an}$ in (12) the constant $3$ becomes $3+o(1)$, but this substitution is inconsequential throughout the analysis (in fact, even increasing the constant, say from 3 to 4, would be inconsequential).

### v3

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Conditioned on the event (21), and given $\tilde{S}$ specified in that event, the SUBSET algorithm succeeds provided that, for all $\ell=1,2,\dotsc,5\eta^{-}k$, and every possible $\tilde{S}^{\prime}$ formed by removing $\ell$ items from $\tilde{S}$ and replacing them with $\ell$ non-defective items, it holds that $\tilde{S}^{\prime}$ explains fewer tests than $\tilde{S}$.
@@
In view of the decoding rule in Algorithm 1, and the fact that $\tilde{S}\in\mathcal{S}_{\rm alg}$, in order to prove the first statement, it suffices to show that the sets $\tilde{S}^{\prime}$ formed in this manner (along with $\tilde{S}$ itself) collectively cover all of $\mathcal{S}_{\rm an}$ (and hence all of $\mathcal{S}_{\rm alg}$, since $\mathcal{S}_{\rm alg}\subseteq\mathcal{S}_{\rm an}$). The range $\ell=1,2,\dotsc,5\eta^{-}k$ suffices because the Hamming distance upon performing $\ell$ swaps is $2\ell$, and $\mathcal{S}_{\rm an}$ constrains the Hamming distance to be at most $10\eta^{-}k$ due to the triangle inequality (see (12)).
@@
    With probability 1o(1)1-o(1), for any $\delta_{2}=\omega(\frac{1}{\sqrt{T}})$, the number of positive tests is at most $\frac{T}{2}(1+\delta_{2})$.
@@
    With probability 1o(1)1-o(1), under the choice $\delta_{2}=\sqrt{\frac{16\eta^{-}k\log\frac{1}{\eta^{-}}}{T}}$, it holds for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ that the number of tests in which at least one item from $\bar{s}$ is present is at least $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$.
@@
We begin with claim (i). Since each test is positive with probability $\frac{1}{2}$ and all tests are mutually independent, the number of positive tests is distributed as $\mathrm{Bin}(T,\frac{1}{2})$. Thus, due to binomial concentration (Appendix -A), it holds that there are at most $\frac{T}{2}(1+\delta_{2})$ positive tests with probability at least $1-\exp(-\frac{\delta_{2}^{2}}{3}T)$, which is 1o(1)1-o(1) when $\delta_{2}=\omega(\frac{1}{\sqrt{T}})$.
@@
We now use (23) along with a union bound to obtain a counterpart that holds simultaneously for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$. The distribution on the number of tests in which $\bar{s}$ is present is the same for any $\bar{s}$, and there are at most ${k\choose(1-\eta^{-})k}={k\choose\eta^{-}k}$ sets in ${\mathcal{S}}_{\mathrm{an}}$ (recall that we implicitly condition on fixed $S=s$), so the probability of any such set being present in less than $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$ tests is upper bounded by
${k\choose\eta^{-}k}\exp\left(-2\eta^{-}k\log\frac{1}{\eta^{-}}\right)$ (24)
@@
Lemma 6 hinges on the two specified assumptions on $\eta^{-}$ and $\theta$; we now proceed to argue that it is safe to assume both of these for the purpose of proving Theorem 2. For $\eta^{-}$, this is because we only replaced the assumed scaling from $\Omega(\cdot)$ to $\Theta(\cdot)$ (and restricted $\lambda$ to $(0,\frac{1}{2})$ rather than $0,\frac{1}{2})$); however, solving SUBSET with a smaller $\eta^{-}$ immediately implies the same for larger $\eta^{-}$, since one can simply remove an arbitrary subset of items from the final estimate. For $\theta$, the restriction $\theta>\frac{1}{3}$ is valid because when $\theta\leq\frac{1}{3}$, a rate of $1$ is achievable even for exact recovery [18, 16], which implies solving SUBSET for arbitrary $\eta^{-}$. Hence, for the rest of the proof, we adopt the stricter assumptions from Lemma 6, in particular noting that $\eta^{-}=o(1)$.
Using part (ii) of Lemma 6 and the fact that $\tilde{S}\in{\mathcal{S}}_{\mathrm{an}}$, we have that $\tilde{S}$ explains at least $T(1-2^{-(1-\eta^{-})(1+o(1))})$ tests with probability 1o(1)1-o(1) (since $\tilde{S}$ explains every test it is present in due to $\tilde{S}\subseteq S$). Hence, and using part (i) of Lemma 6, the number of positive tests not explained by $\tilde{S}$ is at most the following, with probability 1o(1)1-o(1):
@@
[…]
@@
Our proof strategy consists of finding a “worst case distribution” on the number of positive tests the non-defective items in $\mathcal{N}$$_{\ell}$ get placed in, and showing that the probability of $\tilde{S}^{\prime}$ explaining more tests than $\tilde{S}$ under this distribution is small. The probability of at least one of the $\ell$ non-defective items in $\mathcal{N}$$_{\ell}$ being included in any given test is $1-\big(1-\frac{\log 2(1+o(1))}{k}\big)^{\ell}=\frac{\ell\log 2}{k}\big(1+o(1)\big)$, since $\ell=o(k)$. Thus, recalling the high probability upper bound on the number of tests left unexplained by $\tilde{S}$ given in (28), the number of newly explained tests is stochastically upper bounded by the following: [^3] Our analysis focuses on the positive tests that these $\ell$ non-defectives get placed in. Any negative tests they get placed in will only decrease the number of explained tests further, since by Definition 4 any non-defective in a negative test does not explain any tests. However, for our purposes it suffices to naively upper bound this amount of decrease by zero.
@@
We also note the following asymptotic simplification for $2^{\eta^{-}(1+o(1))}-1$ when $\eta^{-}=o(1)$:
@@
With this result established, we are now in a position to show that a union bound over all possible $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$, followed by a second union bound over all the values of $\ell$, still leads to the probability of $\tilde{S}$ not explaining the most tests being small. This is formalized in the following lemma.
@@
Under the setup and notation of Lemma 7, [^4] In particular, we are again conditioning on (21) and (28), and the probability is over the randomness in the placements of non-defective items. for each $\ell\in\{1,2,\dots,5\eta^{-}k\}$, define
${\mathcal{E}}_{\ell}\coloneqq\bigcup_{{\mathcal{D}}_{\ell}\subset\tilde{S}\,:\,\lvert{\mathcal{D}}_{\ell}\rvert=\ell}\bigcup_{{\mathcal{N}}_{\ell}\subset S^{\mathsf{c}}\,:\,\lvert{\mathcal{N}}_{\ell}\rvert=\ell}\left\{T_{\exp}(\tilde{S})\leq T_{\exp}((\tilde{S}\setminus{\mathcal{D}}_{\ell})\cup{\mathcal{N}}_{\ell})\right\}$ (42)
@@
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta^{-}k}{\mathcal{E}}_{\ell}\right)\to 0.$ (43)
@@
We begin by deriving a bound on $\mathbb{P}({\mathcal{E}}_{\ell})$ for fixed $\ell$. For the event corresponding to any single choice of $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$ in (42), the associated probability is upper bounded via Lemma 7. A simple counting argument reveals that there are ${(1-\eta^{-})k\choose\ell}{n-k\choose\ell}$ such pairs, which can be further upper bounded by ${k\choose\ell}{n\choose\ell}$. Hence, the union bound gives
$\mathbb{P}({\mathcal{E}}_{\ell})$ $\leq{k\choose\ell}{n\choose\ell}\mathbb{P}\left(E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}\right)$ (44)
$\leq{k\choose\ell}{n\choose\ell}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (45)
@@
where (46) follows since ${n\choose k}\leq n^{k}$, and (47) follows since $\ell\log k\leq\ell\log n=o(\ell\sqrt{\log n}\cdot\log n)$.
@@
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta^{-}k}{\mathcal{E}}_{\ell}\right)$ $\leq\sum_{\ell=1}^{5\eta^{-}k}\mathbb{P}({\mathcal{E}}_{\ell})$ (48)
$\leq\sum_{\ell=1}^{5\eta^{-}k}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (49)
@@ III-B Proof of Corollary 1
[12, Theorem 1] Fix $\alpha\in(0,1)$ and $\varepsilon>0$, and suppose that $T\geq(1-\alpha)\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Then there exists a group testing algorithm returning an estimate $\hat{S}^{\prime}$ such that $\max\{\lvert S\setminus\hat{S}^{\prime}\rvert,\lvert\hat{S}^{\prime}\setminus S\rvert\leq\alpha k$ with probability 1o(1)1-o(1).
@@
We can use the same idea for the SUBSET problem; since the reasoning is all analogous to that of [12], we omit some details and focus mainly on the differences. We delete a fraction of items arbitrarily close to $\alpha$ and mark them as non-defective. Among those remaining, we apply the approximate recovery algorithm from [18] with $\beta=\eta^{-}$ for some $\eta^{-}=o(1)$ to obtain $\hat{S}^{\prime}$. As noted in [12], a result in [29, Appendix B] shows that this algorithm still succeeds with high probability even when the number of defectives is only known within a factor of $1\pm o(1)$. We then apply the SUBSET algorithm (Algorithm 1) using the reduced ground set and the two-sided estimate $\hat{S}^{\prime}$. While we are unable to directly apply Algorithm 1 and Theorem 2, since these assume that the number of defectives is known, we can easily modify the algorithm and proof to show the result still holds when it is only known to within a factor of $1\pm o(1)$. Specifically, letting $K^{\prime}$ be the (random) number of non-deleted defectives and defining $k^{\prime}\coloneqq\mathbb{E}K^{\prime}=(1-\alpha+\xi)k$, we observe the following:
  • A simple concentration argument reveals that $\underline{k}^{\prime}\leq K^{\prime}\leq\overline{k}^{\prime}$ with probability 1o(1)1-o(1), where $\underline{k}^{\prime}=k^{\prime}(1-o(1))$ and $\overline{k}^{\prime}=k^{\prime}(1+o(1))$.
@@
  • Due to the above modifications, in the definition of ${\mathcal{S}}_{\rm an}$ in (12) the constant $5$ becomes $5+o(1)$, but this substitution is inconsequential throughout the analysis (in fact, even increasing the constant, say from $5$ to $6$, would be inconsequential).